# No Free Lunch from Random Feature Ensembles

### v1

@@ Abstract
Given a budget on total model size, one must decide whether to train a single, large neural network or to combine the predictions of many smaller networks. We study this trade-off for ensembles of random-feature ridge regression models. We prove that when a fixed number of trainable parameters are partitioned among $K$ independently trained models, $K=1$ achieves optimal performance, provided the ridge parameter is optimally tuned. We then derive scaling laws which describe how the test risk of an ensemble of regression models decays with its total size. We identify conditions on the kernel and task eigenstructure under which ensembles can achieve near-optimal scaling laws. Training ensembles of deep convolutional neural networks on CIFAR-10 and a transformer architecture on C4, we find that a single large network outperforms any ensemble of networks with the same total number of parameters, provided the weight decay and feature-learning strength are tuned to their optimal values.
@@ 1 Introduction
A critical consideration in practice is the computational cost associated with ensemble methods. While increasing the number of predictors in an ensemble improves its accuracy (provided each ensemble member is “competent” (Theisen et al., 2023)), each additional model incurs significant computational overhead. Supposing a fixed memory capacity for learned parameters, a more pragmatic comparison is between an ensemble of neural networks and a single large network with the same total parameter count. Indeed, recent studies have called into question the utility of an ensemble of deep networks relative to a single network of comparable total parameter count (Abe et al., 2022; Vyas et al., 2023).
@@
Originally introduced as a fast approximation to Kernel Ridge Regression, random-feature ridge regression (RFRR) Rahimi & Recht (2007) has emerged as a rich “toy model” for deep learning, capturing non-trivial effects of dataset size and network width (Canatar et al., 2021; Atanasov et al., 2023; Mei & Montanari, 2022). While over-fitting effects known as “double-descent” may lead to non-monotonic behavior of the loss in both deep networks and ridge regression (D’Ascoli et al., 2020; Nakkiran, 2019; Nakkiran et al., 2019; Adlam & Pennington, 2020; Lin & Dobriban, 2021), double-descent can be mitigated by optimally tuning the ridge parameter (Nakkiran et al., 2020; Advani et al., 2020; Canatar et al., 2021; Simon et al., 2023). Specifically, Simon et al. showed that in RFRR, test risk decreases monotonically with model size and dataset size when the ridge parameter is optimally chosen.
In this present work, we study the tradeoff between the number of predictors and the size of each predictor in ensembles of RFRR models. We find that with a fixed total parameter budget, minimal error is achieved by a single predictor trained on the full set of features, provided the ridge parameter is optimally tuned. We analyze the trade-off between ensemble size and model size for tasks with a power-law eigenstructure, identifying regimes where near-optimal performance can still be achieved with an ensemble. Finally, we present experiments suggesting that this “no free lunch from ensembles” principle extends to deep feature-learning ensembles when network hyperparameters are finely tuned. The remainder of this work is organized as follows:
In section 2 we review the necessary background on the theory and practice of RFRR and its extension to ensembles.
In section 3, we state an omniscient risk estimate for ensembled RFRR and extend the “bigger is better” theorem of Simon et al. (2023) to the ensembled case.
In Section 4, we prove that, when the total number of features is fixed, the optimal test risk of an ensemble of RFRR models is achieved when the available features are consolidated into a single large model, provided that the ridge parameter is tuned to its optimal value. We confirm these predictions for $\operatorname{ReLU}$ RFRR models on binarized CIFAR-10 and MNIST classification tasks.
In Section 5, we derive scaling laws for the test risk of an ensemble of RFRR models under source and capacity constraints (Cui et al., 2023; Caponnetto & De Vito, 2006; Bordelon et al., 2020; Defilippis et al., 2024), in the width-bottlenecked regime. We identify regimes where near-optimal performance can be achieved using an ensemble of smaller predictors. The derived scaling laws provide a good description of RFRR on the binarized CIFAR-10 and MNIST classification tasks.
In Section 6, we test whether the intuitions provided by random feature models carry over to deep neural networks in computer vision and natural language processing tasks using $\mu P$ parameterization to disentangle the effects of network width and the strength of feature-learning. We find that a single large network outperforms any ensemble of smaller networks, provided that the weight decay and richness are tuned to their optimal values.
@@ 2.1 Random Features and the Kernel Eigenspectrum
In this section, we describe ensembled RFRR, as well as the spectral decomposition of the kernel on which our results rely. This framework is described in (Simon et al., 2023) for the single-predictor case, and reviewed in more rigor in Appendix A.
Kernel Ridge Regression.
In standard kernel ridge regression, the goal is to learn a function $f(\bm{x})$ that maps input features $\bm{x}\in\mathbb{R}^{D}$ to a target value $y\in\mathbb{R}$, given a training set $\mathcal{D}=\{\bm{x}_{p},y_{p}\}_{p=1}^{P}$. The learned function can be expressed in terms of the kernel function $\bm{H}(\bm{x},\bm{x}^{\prime}):\mathbb{R}^{D}\times\mathbb{R}^{D}\mapsto\mathbb{R}$ as:
$f(\bm{x})={\bm{h}}_{x,\mathcal{X}}\left({\bm{H}}_{\mathcal{X}\mathcal{X}}+\lambda\bm{I}\right)^{-1}\bm{y},$ (1)
where $\bm{H}_{\mathcal{X}\mathcal{X}}\in\mathbb{R}^{P\times P}$ is the kernel matrix with entries $\left[\bm{H}_{\mathcal{X}\mathcal{X}}\right]_{pp^{\prime}}=\bm{H}(\bm{x}_{p},\bm{x}_{p^{\prime}})$, and $\bm{h}_{x,\mathcal{X}}=[\bm{H}(\bm{x},\bm{x}_{1}),\dots,\bm{H}(\bm{x},\bm{x}_{P})]$. The vector $\bm{y}\in\mathbb{R}^{P}$ contains the training labels, and $\lambda$ is the ridge parameter.
This procedure can be viewed as performing linear regression in the RKHS defined by the kernel. Specifically, the kernel $\bm{H}(\bm{x},\bm{x}^{\prime})$ can be decomposed into its eigenfunctions $\{\phi_{t}(\bm{x})\}_{t=1}^{\infty}$ and corresponding eigenvalues $\{\eta_{t}\}_{t=1}^{\infty}$:
@@
In this formulation, $f(\bm{x})$ is equivalent to the function learned by linear regression in the infinite-dimensional feature space with dimensions given by $\theta_{t}(\bm{x})\equiv\sqrt{\eta_{t}}\phi_{t}(\bm{x})$. Similarly, we will assume that the target function can be decomposed in this basis as $f_{*}(\bm{x})=\sum_{t}\bar{\bm{w}}_{t}\theta_{t}(\bm{x})$. The training labels are assigned as $y_{p}=f_{*}(\bm{x}_{p})+\epsilon_{p}$ where $\epsilon_{p}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ is drawn i.i.d. for each sample.
Random-Feature Ridge Regression (RFRR).
An approximation of kernel ridge regression may be achieved by mapping the input data into a finite-dimensional feature space, where linear regression is performed. Consider the “featurization” transformation $g:\mathbb{R}^{C}\times R^{D}\mapsto\mathbb{R}$. Define the random features $\bm{\psi}(\bm{x})\in\mathbb{R}^{N}$ by $\left[\bm{\psi}(\bm{x})\right]_{n}=g(\bm{v}_{n},\bm{x})$ for independently drawn $\bm{v}_{n}\sim\mu_{\bm{v}}$. The prediction of the RFRR model is
$f(\bm{x})=\bm{w}^{\top}\bm{\psi}(\bm{x})\,,$ (3)
where $\bm{w}$ is the weight vector learned via ridge regression. The random features model can be interpreted as kernel ridge regression with a stochastic kernel $\hat{\bm{H}}(\bm{x},\bm{x}^{\prime})$, defined as:
$\hat{\bm{H}}(\bm{x},\bm{x}^{\prime})=\frac{1}{N}\sum_{n=1}^{N}g(\bm{v}_{n},\bm{x})g(\bm{v}_{n},\bm{x}^{\prime}).$ (4)
As $N\to\infty$, this stochastic kernel converges to the deterministic kernel $\bm{H}(\bm{x},\bm{x}^{\prime})$. Thus, RFRR provides an approximation to kernel ridge regression that becomes increasingly accurate as the number of random features grows.
@@
Following Simon et al. (2023), we assume that the random features can be replaced by a Gaussian projection from the RKHS associated with the deterministic kernel. Specifically, population risk is well described by the error formula obtained when the random features are replaced by $\bm{\psi}(\bm{x})=\bm{Z}\bm{\theta}(\bm{x})$ where $[\bm{\theta}(\bm{x})]_{t}=\theta_{t}(\bm{x})$, $\bm{Z}\in\mathbb{R}^{N\times H}$ is a random Gaussian matrix with entries $Z_{ij}\sim\mathcal{N}(0,1)$, and $H$ is the (infinite) dimensionality of the RKHS. This assumption is justified rigorously by Defilippis et al. (2024), who provide a multiplicative error bound on the resulting estimate for population risk. The stochastic kernel $\hat{\bm{H}}(\bm{x},\bm{x}^{\prime})$ can be written as the inner product of the random features:
$\hat{\bm{H}}(\bm{x},\bm{x}^{\prime})=\frac{1}{N}\bm{\psi}(\bm{x})^{\top}\bm{\psi}(\bm{x}^{\prime})=\frac{1}{N}\bm{\theta}(\bm{x})^{\top}\bm{Z}^{\top}\bm{Z}\bm{\theta}(\bm{x}^{\prime}).$ (5)
As $N\to\infty$, this stochastic kernel approaches the deterministic kernel $\bm{H}(\bm{x},\bm{x}^{\prime})$.
RFRR Ensembles
Ensembles of RFRR models can be constructed by averaging the predictions made by multiple independently trained RFRR models. For ensemble size $K$, we consider $\bm{\psi}^{k}(\bm{x})\in\mathbb{R}^{N}$, $k=1,\dots,K$ to be the features associated with the $k^{\text{th}}$ ensemble member. The components $\left[\bm{\psi}^{k}(\bm{x})\right]_{n}=g(\bm{v}_{n}^{k},\bm{x})$ for independently drawn $\bm{v}_{n}^{k}\sim\mu_{\bm{v}}$. The ensemble members are trained independently, and then their predictions averaged at test time:
$f_{\text{ens}}(\bm{x})=\frac{1}{K}\sum_{k=1}^{K}f^{k}(\bm{x}),$ (6)
where $f^{k}(\bm{x})=\bm{w}^{k\top}\bm{\psi}^{k}(\bm{x})$ is the prediction of the $k$-th random feature model. Under the assumption of Gaussian universality, we may compute theoretical learning curves by replacing $\bm{\psi}^{k}(\bm{x})=\bm{Z}^{k}\bm{\theta}(\bm{x})\in\mathbb{R}^{N\times H}$, where $\bm{Z}_{1},\dots,\bm{Z}_{K}$ are independently sampled Gaussian random matrices.
@@
The test risk (also known as the generalization error or test error) quantifies the expected error of the learned function on unseen data. In this work, we define the test error as the mean squared error (MSE) between the predicted function $f(\bm{x})$ and the true target function $f(\bm{x})$, averaged over the data distribution $\mu_{\bm{x}}$:
$E_{g}=\mathbb{E}_{\bm{x}\sim\mu_{\bm{x}}}\left[(f(\bm{x})-f_{*}(\bm{x}))^{2}\right]+\sigma_{\epsilon}^{2}.$ (7)
For binary classification problems, we might also consider the clasification error rate on held-out test examples under score-averaging or a majority vote (equations A.6, A.7).

### v2

@@ Abstract
Given a fixed budget for total model size, one must choose between training a single large model or combining the predictions of multiple smaller models. We investigate this trade-off for ensembles of random-feature ridge regression models in both the overparameterized and underparameterized regimes. Using deterministic equivalent risk estimates, we prove that when a fixed number of parameters is distributed among $K$ independently trained models, the ridge-optimized test risk increases with $K$. Consequently, a single large model achieves optimal performance. We then ask when ensembles can achieve near-optimal performance. In the overparameterized regime, we show that, to leading order, the test error depends on ensemble size and model size only through the total feature count, so that overparameterized ensembles consistently achieve near-optimal performance. To understand underparameterized ensembles, we derive scaling laws for the test risk as a function of total parameter count when the ensemble size and parameters per ensemble member are jointly scaled according to a “growth exponent” $\ell$. While the optimal error scaling is always achieved by increasing model size with a fixed ensemble size, our analysis identifies conditions on the kernel and task eigenstructure under which near-optimal scaling laws can be obtained by joint scaling of ensemble size and model size.
Machine Learning, ICML
@@ 1 Introduction
A critical consideration in practice is the computational cost associated with ensemble methods. While increasing the number of predictors in an ensemble improves its accuracy (provided each ensemble member is “competent” (Theisen et al., 2023)), each additional model incurs significant computational overhead. Supposing a fixed memory capacity for learned parameters, a more pragmatic comparison is between an ensemble of neural networks and a single large network with the same total parameter count. Indeed, recent studies have called into question the utility of deep ensembles relative to a single network of comparable total size (Abe et al., 2022; Vyas et al., 2023). However, because ensemble learning allows fully parallelized training and inference, it will likely remain an attractive option in practice. A robust understanding of when an ensemble of small neural networks might approach the performance of a single large neural network is therefore needed.
@@
Originally introduced as a fast approximation to Kernel Ridge Regression, random-feature ridge regression (RFRR) (Rahimi & Recht, 2007) has emerged as a rich “toy model” for deep learning, capturing non-trivial effects of dataset size and network width (Canatar et al., 2021; Atanasov et al., 2023; Mei & Montanari, 2022). In particular, the generalization error of deep networks and RFRR can vary non-monotonically with dataset size and feature count due to over-fitting effects known as “double-descent” (D’Ascoli et al., 2020; Nakkiran, 2019; Nakkiran et al., 2019; Adlam & Pennington, 2020; Lin & Dobriban, 2021), or “multiple descent” (D’Ascoli et al., 2020; Mel & Ganguli, 2021; Meng et al., 2022). However, these overfitting effects can be mitigated with an optimally tuned ridge parameter (Nakkiran et al., 2020; Advani et al., 2020; Canatar et al., 2021; Simon et al., 2023). Specifically, Simon et al. showed that in RFRR, test risk decreases monotonically with model size and dataset size when the ridge parameter is optimally chosen.
In this present work, we study the tradeoff between the number of predictors and the size of each predictor in ensembles of RFRR models. We find that with a fixed budget on total model size (defined as the total number of random features), minimal error is achieved by a single large model, provided the ridge is optimally tuned (theorem 5.1). However, we identify multiple conditions under which ensembles of smaller models can achieve near-optimal performance. In the overparameterized regime, we show that, to leading order, error depends on the ensemble size $K$ and individual model size $N$ only through the total ensemble size $M=KN$ (eq. 6.1), so that overparameterized ensembles consistently achieve near-optimal performance. In the underparameterized regime, we derive scaling laws for the test risk as a function of total ensemble size (eq. 26). The scaling law depends on a newly introduced “growth exponent” $\ell$ which controls the relative growth rates of $K$ and $N$ with total model size $M$ (eq. 25). While optimal error scaling is always achieved by scaling model size $N$ with fixed $K$, we identify conditions on the kernel and task eigenstructure which permit near-optimal error scaling by joint scaling of ensemble size and model size.
@@ 2.1 Random Features and the Kernel Eigenspectrum
In this section, we describe ensembled RFRR, as well as the spectral decomposition of the kernel on which our results rely. This framework is described in (Simon et al., 2023) for the single-predictor case, and is reviewed in more rigor in Appendix A. Here we will consider a supervised learning task, where the goal is to learn an estimator $\hat{f}(\bm{x})$ that maps input features $\bm{x}\in\mathbb{R}^{D}$ to a target value $y\in\mathbb{R}$, given a training set $\mathcal{D}=\{\bm{x}_{p},y_{p}\}_{p=1}^{P}$. The training labels are assigned as $y_{p}=f_{*}(\bm{x}_{p})+\epsilon_{p}$ for some ground truth function $f_{*}$, and where $\epsilon_{p}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ is drawn i.i.d. for each sample.
Random-Feature Ridge Regression (RFRR) Ensembles.
In a Random-Feature Ridge Regression (RFRR) Ensemble of size $K$, the data is first transformed into $K$ separate finite-dimensional feature-spaces. Linear regression is performed on each of these spaces independently, and the resulting predictions are averaged at test time. Throughout this paper, we will use upper indices $k=1,\dots,K$ to represent the ensemble index. $K$ sets of random features $\bm{\psi}^{k}(\bm{x})\in\mathbb{R}^{N}$ are constructed as $\left[\bm{\psi}^{k}(\bm{x})\right]_{n}=g(\bm{v}_{n}^{k},\bm{x})$. Here, for each $n=1,\dots,N$ and $k=1,\dots,K$, $\bm{v}_{n}^{k}$ is a random parameter vector sampled independently as $\bm{v}_{n}^{k}\sim\mu_{\bm{v}}$ for measure $\mu_{\bm{v}}$ on $\mathbb{R}^{C}$. The function $g:\mathbb{R}^{C}\times\mathbb{R}^{D}\mapsto\mathbb{R}$ is a “featurization transformation,” often taking the form $g(\bm{v},\bm{x})=\varphi(\bm{v}^{\top}\bm{x})$ for some nonlinear activation function $\varphi(\cdot)$. The prediction of the $k^{\text{th}}$ RFRR model is then given by
$\hat{f}^{k}(\bm{x})=\frac{\hat{\bm{w}}^{k\top}\bm{\psi}^{k}(\bm{x})}{\sqrt{N}}\,$, (1)
where $\hat{\bm{w}}^{k}$ is the weight vector learned via ridge regression:
[…]
$=\left(\frac{1}{N}{\bm{\Psi}^{k}}^{\top}\bm{\Psi}^{k}+\lambda\mathbf{I}\right)^{-1}\frac{{\bm{\Psi}^{k}}^{\top}\bm{y}}{\sqrt{N}}$ (3)
where $\bm{\Psi}^{k}\in\mathbb{R}^{N\times P}$ has columns $\bm{\psi}^{k}(\bm{x}_{1}),\cdots,\bm{\psi}^{k}(\bm{x}_{P})]$ and the vector $\bm{y}\in\mathbb{R}^{P}$ has $\bm{y}]_{p}=y_{p}$. Finally, the ensembled predictor is constructed by averaging the predictions of each ensemble member $k=1,\dots,K$:
$\hat{f}_{\text{ens}}(\bm{x})=\frac{1}{K}\sum_{k=1}^{K}\hat{f}^{k}(\bm{x}),$ (4)
The Kernel Limit
Each of the learned functions $\hat{f}^{k}(\bm{x})$ may equivalently be expressed as the kernel ridge regression predictor:
$\hat{f}^{k}(\bm{x})=\hat{\bm{h}}^{k}_{x,\mathcal{X}}\left(\hat{\bm{H}}^{k}_{\mathcal{X}\mathcal{X}}+\lambda\bm{I}_{N}\right)^{-1}\bm{y}$ (5)
Where the matrix $\hat{\bm{H}}^{k}_{\mathcal{X}\mathcal{X}}]_{pp^{\prime}}=\hat{\bm{H}}^{k}(\bm{x}_{p},\bm{x}_{p^{\prime}})$ and the vector $\hat{\bm{h}}^{k}_{x,\mathcal{X}}]_{p}=\hat{\bm{H}}^{k}(\bm{x},\bm{x_{p}})$, for the stochastic finite-feature kernels:
$\hat{\bm{H}}^{k}(\bm{x},\bm{x}^{\prime})=\frac{1}{N}\sum_{n=1}^{N}g(\bm{v}_{n}^{k},\bm{x})g(\bm{v}_{n}^{k},\bm{x}^{\prime})$ (6)
The fluctuations in this stochastic kernel vary due to the random realization of the weights $\{\bm{v}_{n}^{k}\}_{n=1}^{N}$ across ensemble members $k=1,\dots,K$. As $N\to\infty$, these stochastic kernels converge to a common deterministic limiting kernel:
$\hat{\bm{H}}^{k}(\bm{x},\bm{x}^{\prime})\xrightarrow{N\to\infty}\bm{H}(\bm{x},\bm{x}^{\prime})=\mathbb{E}_{\bm{v}\sim\mu_{\bm{v}}}\left[g(\bm{v},\bm{x})g(\bm{v},\bm{x}^{\prime})\right]$
In this “kernel limit”, the learned functions $\hat{f}^{k}(\bm{x})$ will be independent of the ensemble index $k$, so that we can define $\hat{f}_{\infty}(\bm{x})\equiv\hat{f}_{ens}(\bm{x})=\hat{f}^{k}(\bm{x})\,\forall\,k$, with
$\hat{f}_{\infty}(\bm{x})={\bm{h}}_{x,\mathcal{X}}\left({\bm{H}}_{\mathcal{X}\mathcal{X}}+\lambda\bm{I}\right)^{-1}\bm{y}\,,\quad(N\to\infty).$ (7)
where $\bm{H}_{\mathcal{X}\mathcal{X}}\in\mathbb{R}^{P\times P}$ is the kernel matrix with entries $\left[\bm{H}_{\mathcal{X}\mathcal{X}}\right]_{pp^{\prime}}=\bm{H}(\bm{x}_{p},\bm{x}_{p^{\prime}})$, and $\bm{h}_{x,\mathcal{X}}=[\bm{H}(\bm{x},\bm{x}_{1}),\dots,\bm{H}(\bm{x},\bm{x}_{P})]$. The limiting kernel $\bm{H}(\bm{x},\bm{x}^{\prime})$ can be decomposed into its eigenfunctions $\{\phi_{t}(\bm{x})\}_{t=1}^{\infty}$ and corresponding eigenvalues $\{\eta_{t}\}_{t=1}^{\infty}$:
@@
In this formulation, $\hat{f}(\bm{x})$ is equivalent to the function learned by linear regression in the infinite-dimensional features with components given by $\theta_{t}(\bm{x})\equiv\sqrt{\eta_{t}}\phi_{t}(\bm{x})$. We will assume that the target function can be decomposed in this basis as $f_{*}(\bm{x})=\sum_{t}\bar{\bm{w}}_{t}\theta_{t}(\bm{x})$. In the following sections, we will describe an error formula for RFRR ensembles which depends on these limiting kernel eigenvalues $\{\eta_{t}\}_{t=1}^{\infty}$ and target weights $\{\bm{w}_{t}\}_{t=1}^{\infty}$, $K$, NN, and $P$, which provide a complete statistical description of the learning problem.
@@
Following (Simon et al., 2023), we assume that the random features can be replaced by a Gaussian projection from the RKHS associated with the limiting kernel. Specifically, population risk is well described by the error formula obtained when the random features are replaced by $\bm{\psi}^{k}(\bm{x})=\bm{Z}^{k}\bm{\theta}(\bm{x})$ where $\bm{\theta}(\bm{x})]_{t}=\theta_{t}(\bm{x})$, $\bm{Z}^{k}\in\mathbb{R}^{N\times H}$ is a random Gaussian matrix with entries $Z^{k}_{ij}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$, and $H$ is the (infinite) dimensionality of the RKHS. Equivalently, we may approximate the stochastic kernels $\hat{\bm{H}}^{k}(\bm{x},\bm{x}^{\prime})$ as the inner products of linear random features:
$\hat{\bm{H}}^{k}(\bm{x},\bm{x}^{\prime})=\frac{1}{N}\bm{\psi}^{k}(\bm{x})^{\top}\bm{\psi}^{k}(\bm{x}^{\prime})=\frac{1}{N}\bm{\theta}(\bm{x})^{\top}\bm{Z}^{k\top}\bm{Z}^{k}\bm{\theta}(\bm{x}^{\prime}).$ (9)
Under this Gaussian universality assumption, the RFRR setting is equivalent to the linear random-features setting studied in ref’s. (Maloney et al., 2022; Atanasov et al., 2024, 2022), with the limiting kernel eigenspectrum playing the role of the data covariance.
The Gaussian universality assumption has been justified rigorously in the case of a single ensemble member ($K=1$) by Defilippis et al., who provide a multiplicative error bound on the resulting estimate for population risk (see section 3 and Appendix B).
@@
The test risk (also known as the generalization error or test error) quantifies the expected error of the learned function on unseen data. In this work, we define the test error as the mean squared error (MSE) between the predicted function $f_{\text{ens}}(\bm{x})$ and the true target function $f_{*}(\bm{x})$, averaged over the data distribution $\mu_{\bm{x}}$:
$\mathcal{E}_{g}^{K}=\mathbb{E}_{\bm{x}\sim\mu_{\bm{x}}}\left[(\hat{f}_{\text{ens}}(\bm{x})-f_{*}(\bm{x}))^{2}\right]+\sigma_{\epsilon}^{2}.$ (10)
For binary classification problems, we might also consider the classification error rate on held-out test examples under score-averaging or a majority vote (equations A.6, A.7).